A 2.00(g) sample of a mixture of CaCO(s) and BaO(s) is placed in a 1.00 L container containing CO2(g) at a pressure of 785 torr and a T = 525 C. The CO2(g) reacts with the solids forming calcium carbonate and barium carbonate to completion. The pressure of the CO2(g) after the reaction is complete is 964 Pa. Write the chemical reactions and calculate the mole fraction X of CaO and BaO. Please show all work.
Answer #1

Ans:

Mass of the mixture = 2 g

Let ‘x’ g be the mass of CaO in the mixture

Then mass of BaO in the mixture = (2-x)g

Molar mass of CaO = 56.1 g/mol

Molar mass of BaO = 153.3 g/mol

No. of moles of CaO taken = (x/56.1) mol

No. of moles of BaO taken = {(2-x)/153.3} mol

Volume of the container, V = 1 L = 10-3 m3

Temperature, T = 525 0C = 798 K

Pressure, P = 785 torr = 104658 Pa (1 torr = 133.322 Pa)

n = no. of moles of gas

R = gas constant = 8.314 m3⋅Pa⋅K−1⋅mol−1

According to Ideal gas equation:

PV = nRT & n = PV/RT

n = {(104658 Pa x 10-3 m3)/( 8.314 m3⋅Pa⋅K−1⋅mol−1 x 798 K)} = 0.015775 mol

No. of moles of CO2 taken initially = 0.015775 mol

==> Pressure of CO2 after the reaction = 964 Pa

According to Ideal gas equation:

PV = nRT & n = PV/RT

n = {(964 Pa x 10-3 m3)/( 8.314 m3⋅Pa⋅K−1⋅mol−1 x 798 K)} = 0.000145 mol

No of moles of CO2 reacted = (0.015775 mol - 0.000145 mol) = 0.01563 mol

Reaction of CaO & BaO with CO2:

CaO + CO2 --> CaCO3 (1 mol of CaO reacts with 1 mol of CO2 giving 1 mol of CaCO3)

BaO + CO2 --> BaCO3 (1 mol of BaO reacts with 1 mol of CO2 giving 1 mol of BaCO3)

Since CaO & BaO reacts with CO2 in 1:1 ratio, we can tell (x/56.1) mol of CaO reacts with (x/56.1) mol of CO2 and {(2-x)/153.3} mol of BaO reacts with {(2-x)/153.3} mol of CO2.

So, (x/56.1)mol + 1{(2-x)/153.3} mol = 0.01563 mol

Solving for ‘x’ we get, x = 0.2286 g

No. of moles of CaO = x/56.1 = 0.2286/56.1 = 0.0041 mol (n1)

No. of moles of BaO = (2-x)/153.3 = (2-0.2286)/153.3 = 0.1156 mol (n2)

Mole fraction of CaO = n1/(n1+n2) = 0.0041/(0.0041+0.1156) = 0.0343

Mole fraction of BaO = (1 – mole fraction of CaO) = (1- 0.0343) = 0.9657